# Unified Privacy Guarantees for Decentralized Learning via Matrix Factorization

### v1

@@ Appendix A SotA DP-DL Approaches as Matrix Factorization Mechanisms
In this section we prove Theorem 5, by looking at algorithms one by one. But first we explain the strategy used to get the factorization, and we present a useful Lemma.
@@ A.1 Path to Matrix Factorization Mechanisms
For each setting the proof of Theorem 5 follows a two-steps strategy:
@@ A.2 Proof of Theorem 5 for DP-D-SGD
With LDP, the attacker has only access to messages, therefore no reduction is needed and we analyze directly the factorization of the messages. We have $C_{\text{DP--SGD}}=I_{nT}$, yielding $A_{\text{DP-D-SGD}}=B_{\text{DP-D-SGD}}={\mathbf{W}}_{T}$.
@@ A.3 Proof of Theorem 5 for ANTI-PGD
In this section, we make explicit the factorization of Theorem 5 for the AntiPGD algorithm, which was studied in Koloskova et al. [2023].
@@
Following the factorization found in a centralized setting [Koloskova et al., 2023], we have $C_{\text{{Anti-PGD}}}=1_{t\geq t^{\prime}}\otimes I_{n}$, $A_{\text{ANTI-PGD}}={\mathbf{W}}_{T}$ and:
@@ A.4 Proof of Theorem 5 for Muffliato-SGD
In this section, we make explicit the factorization of Theorem 5 for the Muffliato algorithm, which was studied in Cyffers et al. [2022].
@@
Muffliato-SGD [Cyffers et al., 2022] requires a different communication workload than ${\mathbf{W}}_{T}$. That is because they consider multiple communication rounds before computing a new gradient. Formally, they define $K$ the number of communication steps between two gradients, and they perform $T$ iterations. This means that the set of messages shared is:
@@
PNDP considers a set $\cal A$ of attacker nodes. To ease the proof, without loss of generality, we reorder G and Z such that their first lines corresponds to the attackers gradients and noise.
@@ A.5 Proof of Theorem 5 for DECOR
In this section, we make explicit the factorization of Theorem 5 for the DECOR algorithm, which was studied in Allouah et al. [2024].
@@
With DECOR, each node knows some of the noises of its neighbors. Therefore Sec-LDP is more appropriate to describe the knowledge gathered by some attackers.
@@
Let consider a set $\cal A$ of attacker nodes. To ease the proof, without loss of generality, we reorder G such that its first lines corresponds to the attackers gradients. For Z, we first gather the local noise of the attackers, then other local noises, and finally the ”secrets” noise.
@@ A.6 Proof of Theorem 5 for Zip-DL
In this section, we make explicit the factorization of Theorem 5 for the Zip-DL algorithm, which was introduced in Biswas et al. [2025].
@@
Zip-DL is already expressed under a matrix form [Biswas et al., 2025]: by introducing $n^{2}$ virtual nodes, and translating express the gossip matrix W{W} under this virtual equivalent as $\widetilde{W}\in\mathbb{R}^{n^{2}\times n^{2}}$ to encodes messages. Formally, we have $\widetilde{W}[ni+j+1,nj+i+1]={W}[i,j]$. Furthermore, considering an averaging matrix $\widetilde{M}$, and a correlation matrix $C_{\text{Zip-DL}}^{\dagger}$ (coined $\hat{C}$ in Biswas et al. [2025]), Zip-DL can be expressed as:
@@
For the ease of notation, we assume here that for any node $i$, $s_{i}=c$. The generalization is straight-forward. To ease the proof, without loss of generality, we reorder $G$ and $Z$ such that their first lines correspond to the attackers gradients and noise.
@@ Appendix B Proof of Theorem 7
We denote $\mathcal{M}\equiv\mathcal{M}^{\prime}$ two mechanism that have the same $\mu$-GDP properties. In this proof, this will either be because they are distributionally equivalent, or because they are an invertible operation away from each other (meaning they have the same $\mu$-GDP properties by post-processing).
@@
Consider an $LQ$-decomposition of $B^{\prime}$ : since $\operatorname{rank}(B^{\prime})=\operatorname{rank}(B)\leq m$ we can find a decomposition $B^{\prime}=\begin{bmatrix}L&0\end{bmatrix}Q$ such that $L\in\mathbb{R}^{\operatorname{rank}(B^{\prime})\times\operatorname{rank}(B^{\prime})}$ is lower triangular and $Q\in\mathbb{R}^{m\times m}$ is an orthonormal matrix Trefethen [1997]. Then, $\mathcal{M}_{1}(G)=\begin{bmatrix}L&0\end{bmatrix}(QCG+QZ)$ for any $G$. Since $Q$ is an orthonormal matrix, this is distributionally equivalent to $\mathcal{M}_{2}$.
@@
$L$ in $\mathcal{M}_{2}$ has been constructed to be square and lower triangular matrix of full rank (as $B^{\prime}$ is of full-rank). Thus, $L$ is invertible. This means, by post-processing using an invertible operator $L$, that the privacy guarantees of $\mathcal{M}_{2}$ are the one of $\begin{bmatrix}I_{\operatorname{rank}(B^{\prime})}&0\end{bmatrix}(QCG+Z)$. It suffices to name $\tilde{Q}=\begin{bmatrix}I_{\operatorname{rank}(B^{\prime})}&0\end{bmatrix}Q$ and $\tilde{Z}=\begin{bmatrix}I_{\operatorname{rank}(B^{\prime})}&0\end{bmatrix}Z$ to see this corresponds to $\mathcal{M}_{3}$.
@@
Then, applying Lemma 3.15 of Pillutla et al. [2025] allows us to derive the following bound:
@@
where Equation 28 corresponds to Part (a) of Lemma 3.15 in Pillutla et al. [2025] (generalization to batch releases). We also have equality when the $\left(\left(\tilde{Q}C\right)^{\top}\tilde{Q}C\right)[t,\tau]>0$ for all $t,\tau\in\pi$, for all $\pi\in\Pi$).
@@ C.1 Housing privacy-utility tradeoff
We also introduce D-MF as a baseline, an adaptation of the local correlation introduced by Choquette-Choo et al. 2023 where each node locally performs the correlation derived for centralized and federated settings. We report and compare the result of this baseline in Figure 4.
Smaller graphs only yield better privacy/utility tradeoffs only because data is partitioned in bigger batches on smaller graphs, as there is more data for each node, netting better privacy properties.
(a) Ego, $\mu=0.5$
(b) Peertube, $\mu=1$
Figure 5: Evolution of the test loss on various $\mu$ and graph settings.
Figure 5 presents the evolution of the test loss for specific privacy guarantees $\mu$ and two graphs — Ego and Peertube. In all cases, we observe an improvement over all baselines.
@@ Appendix D Proof of Lemma 10
As highlighted in Loan [2000], the commutation matrix $\mathbf{K}^{(n,T)}$ (also called the vec-permutation matrix Henderson & Searle [1981]) is a way to change the representation from stacking through time the observations of the global system to stacking locally the observations through time for each node. In other words, for any matrix $X\in\mathbb{R}^{n\times n},Y\in\mathbb{R}^{T\times T}$, we have:
@@
It suffices to focus on the norm, since the sensitivity term remains unchanged. Let $A=\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\mathbf{K}_{n,T}$. First, note that $\mathbf{K}_{T,n}$ is a permutation matrix, so we have:
$\left\lVert\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\left(C_{local}^{\dagger}\otimes I_{n}\right)\right\rVert_{F}^{2}$ $=\left\lVert\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\left(C_{local}^{\dagger}\otimes I_{n}\right)\mathbf{K}_{T,n}\right\rVert_{F}^{2}$
$=\left\lVert\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\mathbf{K}_{n,T}\left(I_{n}\otimes C_{local}^{\dagger}\right)\right\rVert_{F}^{2}$

### v2

@@ Appendix A SotA DP-DL Approaches as Matrix Factorization Mechanisms
In this section, we prove Theorem 6 by looking at algorithms one by one. But first, we explain the strategy used to get the factorization, and we present a useful Lemma.
@@ A.1 Path to Matrix Factorization Mechanisms
For each setting, the proof of Theorem 6 follows a two-step strategy:
@@ A.2 Proof of Theorem 6 for DP-D-SGD
With LDP, the attacker has only access to messages, therefore, no reduction is needed, and we analyze the factorization of the messages directly. We have $C_{\text{DP--SGD}}=I_{nT}$, yielding $A_{\text{DP-D-SGD}}=B_{\text{DP-D-SGD}}={\mathbf{W}}_{T}$.
@@ A.3 Proof of Theorem 6 for ANTI-PGD
In this section, we make explicit the factorization of Theorem 6 for the AntiPGD algorithm, which was studied in Koloskova et al. (2023).
@@
Following the factorization found in a centralized setting (Koloskova et al., 2023), we have $C_{\text{{Anti-PGD}}}=1_{t\geq t^{\prime}}\otimes I_{n}$, $A_{\text{ANTI-PGD}}={\mathbf{W}}_{T}$ and:
@@ A.4 Proof of Theorem 6 for Muffliato-SGD
In this section, we make explicit the factorization of Theorem 6 for the Muffliato algorithm, which was studied in Cyffers et al. (2022).
@@
Muffliato-SGD (Cyffers et al., 2022) requires a different communication workload than ${\mathbf{W}}_{T}$. That is because they consider multiple communication rounds before computing a new gradient. Formally, they define $K$ the number of communication steps between two gradients, and they perform $T$ iterations. This means that the set of messages shared is:
@@ A.5 Proof of Theorem 6 for DECOR
In this section, we make explicit the factorization of Theorem 6 for the DECOR algorithm, which was studied in Allouah et al. (2024).
@@
With DECOR, each node knows some of the noises of its neighbors. Therefore, Sec-LDP is more appropriate to describe the knowledge gathered by some attackers.
@@
Let us consider a set $\cal A$ of attacker nodes. To ease the proof, without loss of generality, we reorder G such that its first lines correspond to the attackers gradients. For Z, we first gather the local noise of the attackers, then other local noises, and finally the ”secrets” noise.
@@ A.6 Proof of Theorem 6 for Zip-DL
In this section, we make explicit the factorization of Theorem 6 for the Zip-DL algorithm, which was introduced in Biswas et al. (2025).
@@
Zip-DL is already expressed under a matrix form (Biswas et al., 2025): by introducing $n^{2}$ virtual nodes, and translating to express the gossip matrix W{W} under this virtual equivalent as $\widetilde{W}\in\mathbb{R}^{n^{2}\times n^{2}}$ to encode messages. Formally, we have $\widetilde{W}[ni+j+1,nj+i+1]={W}[i,j]$. Furthermore, considering an averaging matrix $\widetilde{M}$, and a correlation matrix $C_{\text{Zip-DL}}^{\dagger}$ (coined $\hat{C}$ in Biswas et al. (2025)), Zip-DL can be expressed as:
@@
For the ease of notation, we assume here that for any node $i$, $s_{i}=c$. The generalization is straightforward. To ease the proof, without loss of generality, we reorder $G$ and $Z$ such that their first lines correspond to the attackers gradients and noise.
@@
A.7 Extending notations to time-varying gossip
We extend the notations introduced in Section 4.1 to time-varying graphs encoded as a sequence of gossip matrices ${W}_{t}$, for $t\in\left\llbracket 1,T\right\rrbracket$. In this more general context, the workload matrix ${\mathbf{W}}_{T}$ becomes:
${\mathbf{W}}_{T}=\begin{bmatrix}I_{n}&0&0&\ldots&0\\ {W}_{1}&I_{n}&0&\ldots&0\\ {W}_{2}{W}_{1}&{W}_{2}&I_{n}&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \Pi_{t=T}^{1}{W}_{t}&\Pi_{t=T-1}^{2}{W}_{t}&\Pi_{t=T-1}^{3}{W}_{t}&\ldots&I_{n}\end{bmatrix}$
This new definition of ${\mathbf{W}}_{T}$ encodes how messages spread through time using the sequence of gossip matrices $({W}_{t})_{t}$. The formalization of the MF-D-SGD algorithm (initially presented in Equation 5) similarly needs adapting by replacing $I_{T}\otimes{W}$ by a block diagonal matrix of all ${W}_{t}$:
$\theta\;=\;\begin{bmatrix}{W}_{1}&0&\ldots&0\\ 0&{W}_{2}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&{W}_{T}\\ \end{bmatrix}\left(M\theta_{0}-\eta\,{\mathbf{W}}_{T}\bigl(G+C^{\dagger}Z\bigr)\right),$ (23)
@@ Appendix B Proof of Theorem 8
We denote $\mathcal{M}\equiv\mathcal{M}^{\prime}$ two mechanisms that have the same $\mu$-GDP properties. In this proof, this will either be because they are distributionally equivalent, or because they are an invertible operation away from each other (meaning they have the same $\mu$-GDP properties by post-processing).
@@
Consider an $LQ$-decomposition of $B^{\prime}$ : since $\operatorname{rank}(B^{\prime})=\operatorname{rank}(B)\leq m$ we can find a decomposition $B^{\prime}=\begin{bmatrix}L&0\end{bmatrix}Q$ such that $L\in\mathbb{R}^{\operatorname{rank}(B^{\prime})\times\operatorname{rank}(B^{\prime})}$ is lower triangular and $Q\in\mathbb{R}^{m\times m}$ is an orthonormal matrix Trefethen (1997). Then, $\mathcal{M}_{1}(G)=\begin{bmatrix}L&0\end{bmatrix}(QCG+QZ)$ for any $G$. Since $Q$ is an orthonormal matrix, this is distributionally equivalent to $\mathcal{M}_{2}$.
@@
$L$ in $\mathcal{M}_{2}$ has been constructed to be square and lower triangular matrix of full rank (as $B^{\prime}$ is of full rank). Thus, $L$ is invertible. This means, by post-processing using an invertible operator $L$, that the privacy guarantees of $\mathcal{M}_{2}$ are the one of $\begin{bmatrix}I_{\operatorname{rank}(B^{\prime})}&0\end{bmatrix}(QCG+Z)$. It suffices to name $\tilde{Q}=\begin{bmatrix}I_{\operatorname{rank}(B^{\prime})}&0\end{bmatrix}Q$ and $\tilde{Z}=\begin{bmatrix}I_{\operatorname{rank}(B^{\prime})}&0\end{bmatrix}Z$ to see this corresponds to $\mathcal{M}_{3}$.
@@
Then, applying Lemma 3.15 of Pillutla et al. (2025) allows us to derive the following bound:
@@
where Equation 29 corresponds to Part (a) of Lemma 3.15 in Pillutla et al. (2025) (generalization to batch releases). We also have equality when the $\left(\left(\tilde{Q}C\right)^{\top}\tilde{Q}C\right)[t,\tau]>0$ for all $t,\tau\in\pi$, for all $\pi\in\Pi$).
@@ C.1 Housing privacy-utility tradeoff
Smaller graphs only yield better privacy/utility tradeoffs because data is partitioned into bigger batches on smaller graphs, as there is more data for each node, netting better privacy properties.
C.2 Femnist privacy-utility tradeoff
Florentine Families
Facebook Ego graph
Figure 6: Privacy-utility tradeoff on the FEMNIST dataset for various graphs.
We also include Figure 6 to present the privacy-utility tradeoffs for the FEMNIST dataset on the Florentine families graph and the Facebook Ego graph. In both cases, better test accuracy is consistently reached by Mafalda-SGD for different values of $\epsilon$.
@@ Appendix D Proof of Lemma 12
As highlighted in Loan (2000), the commutation matrix $\mathbf{K}^{(n,T)}$ (also called the vec-permutation matrix Henderson and Searle (1981)) is a way to change the representation from stacking through time the observations of the global system to stacking the observations locally of each node through time. In other words, for any matrix $X\in\mathbb{R}^{n\times n},Y\in\mathbb{R}^{T\times T}$, we have:
@@
It suffices to focus on the norm, since the sensitivity term remains unchanged. Let $A=\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\mathbf{K}^{(n,T)}$. First, note that $\mathbf{K}^{(T,n)}$ is a permutation matrix, so we have:
$\left\lVert\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\left(C_{local}^{\dagger}\otimes I_{n}\right)\right\rVert_{F}^{2}$ $=\left\lVert\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\left(C_{local}^{\dagger}\otimes I_{n}\right)\mathbf{K}^{(T,n)}\right\rVert_{F}^{2}$
$=\left\lVert\left(I_{T}\otimes W\right){\mathbf{W}}_{T}\mathbf{K}^{(n,T)}\left(I_{n}\otimes C_{local}^{\dagger}\right)\right\rVert_{F}^{2}$